# Ctrl-X: Controlling Structure and Appearance for Text-To-Image Generation Without Guidance

### v1

@@ Appendix B Extension to prompt-driven controllable generation
Ctrl-X also supports prompt-driven conditional generation, where it generates an output image complying with the given text prompt while aligning with the structure from the structure image, as shown in Figures 4 and 6. Inspired by FreeControl [24], instead of a given $\mathbf{I}^{\mathrm{a}}$, Ctrl-X can jointly generate $\mathbf{I}^{\mathrm{a}}$ based on the text prompt alongside $\mathbf{I}^{\mathrm{o}}$, where we obtain $\mathbf{x}^{\mathrm{a}}_{t-1}$ via denoising with Equation 2 from $\mathbf{x}^{\mathrm{a}}_{t}$ without control.
@@
Dataset.    Our controllable generation dataset comprises of $175$ diverse image-prompt pairs with the same (structure) images as Section 5.1. It consists of $71\%$ ControlNet-supported conditions and $29\%$ new conditions. We use the same hand-annotated structure prompts and hand-create output prompts with inspiration from Plug-and-Play’s datasets [36]. See more details in Appendix D.
@@
Qualitative results.    As shown in Figures 4 and 10, Ctrl-X generates high-quality images with great structure preservation and close prompt alignment. Our method can extract structure information from a wide range condition types and produces results of diverse modalities based on the prompt.
@@
Method ControlNet-supported New condition
Self-sim $\downarrow$ CLIP score $\uparrow$ LPIPS $\uparrow$ Self-sim $\downarrow$ CLIP score $\uparrow$ LPIPS $\uparrow$
ControlNet [44] 0.126 0.298 0.657 0.092 0.302 0.507
T2I-Adapter [25] 0.096 0.303 0.504 0.068 0.302 0.415
SDEdit [23] 0.102 0.300 0.366 0.096 0.309 0.373
Prompt-to-Prompt [11] 0.100 0.276 0.370 0.097 0.287 0.357
Plug-and-Play [36] 0.056 0.282 0.272 0.050 0.292 0.301
InfEdit [41] 0.117 0.314 0.523 0.102 0.311 0.442
FreeControl [24] 0.108 0.340 0.557 0.104 0.339 0.492
Ctrl-X (ours) 0.134 0.322 0.635 0.135 0.326 0.590
@@
Appendix C Additional video results
@@
Figure 11: Additional text-to-video (T2V) model results.    Ctrl-X can be directly applied to T2V models for controllable video structure and appearance control, with AnimateDiff [9] with Realistic Vision v5.1 [32] and LaVie [39] here as examples.
@@
We present additional results of our method directly applied to text-to-video (T2V) diffusion models in Figure 11, namely AnimateDiff [9] with base model Realistic Vision v5.1 [32] and LaVie [39].
Appendix D Dataset details
All academic datasets which we use are cited here [3, 10, 43, 24, 48, 22, 19, 36, 16]. We will release our dataset alongside code release.
@@ Appendix D Dataset details
Our dataset consists of 177 $1024\times 1024$ images divided into 16 types and across 7 categories. We split the images into condition images (67 images: “canny edge map”, “metadrive”, “3d mesh”, “3d humanoid”, “depth map”, “human pose image”, “point cloud”, “sketch”, “line drawing”, “HED edge drawing”, “normal map”, and “segmentation mask”) and natural images (110 images: “photo”, “painting”, “cartoon” and “birds eye view”), with the the largest type being “photo” (83 images). The condition images are further divided into two groups in our paper: ControlNet-supported conditions (“canny edge map”, “depth map”, “human pose image”, “line drawing”, “HED edge drawing”, “normal map”, and “segmentation mask”) and in-the-wild conditions (“metadrive”, “3D mesh”, “3D humanoid”, “point cloud”, and “sketch”). All of our images fall into one of seven categories: “animals” (52 images), “buildings” (11 images), “humans” (28 images), “objects” (29 images), “rooms” (24 images), “scenes” (22 images) and “vehicles” (11 images). About two thirds of the images come from the Web, while the remaining third is generated using SDXL 1.0 [27] or converted from natural images using Controlnet Annotators packaged in controlnet-aux [44]. For each of these images, we hand annotate them with a text prompt and other metadata (e.g. type). Then, these images, promtps, and metadata are combined to form the structure and appearance control dataset and conditional generation dataset, detailed below.
@@
In practice, when running Ctrl-X, we simply set the appearance prompt to be the same as the output prompt instead of our hand-annotated appearance prompt. We found little difference between the two approaches.

### v2

@@ Appendix C Extension to prompt-driven controllable generation
Ctrl-X also supports prompt-driven conditional generation, where it generates an output image complying with the given text prompt while aligning with the structure from the structure image, as shown in Figures 4 and 7. Inspired by FreeControl [24], instead of a given $\mathbf{I}^{\mathrm{a}}$, Ctrl-X can jointly generate $\mathbf{I}^{\mathrm{a}}$ based on the text prompt alongside $\mathbf{I}^{\mathrm{o}}$, where we obtain $\mathbf{x}^{\mathrm{a}}_{t-1}$ via denoising with Equation 2 from $\mathbf{x}^{\mathrm{a}}_{t}$ without control.
@@
Dataset.    Our controllable generation dataset comprises of $175$ diverse image-prompt pairs with the same (structure) images as Section 5.1. It consists of $71\%$ ControlNet-supported conditions and $29\%$ new conditions. We use the same hand-annotated structure prompts and hand-create output prompts with inspiration from Plug-and-Play’s datasets [36]. See more details in Appendix E.
@@
Qualitative results.    As shown in Figures 4 and 13, Ctrl-X generates high-quality images with great structure preservation and close prompt alignment. Our method can extract structure information from a wide range condition types and produces results of diverse modalities based on the prompt.
@@
Method Training ControlNet-supported New condition Self-sim $\downarrow$ CLIP score $\uparrow$ LPIPS $\uparrow$ Self-sim $\downarrow$ CLIP score $\uparrow$ LPIPS $\uparrow$ ControlNet [44] 0.126 0.298 0.657 0.092 0.302 0.507 T2I-Adapter [25] 0.096 0.303 0.504 0.068 0.302 0.415 SDEdit [23] 0.102 0.300 0.366 0.096 0.309 0.373 Prompt-to-Prompt [11] 0.100 0.276 0.370 0.097 0.287 0.357 Plug-and-Play [36] 0.056 0.282 0.272 0.050 0.292 0.301 InfEdit [41] 0.117 0.314 0.523 0.102 0.311 0.442 FreeControl [24] 0.108 0.340 0.557 0.104 0.339 0.492 Ctrl-X (ours) 0.134 0.322 0.635 0.135 0.326 0.590
@@
Appendix D Additional results
Figure 14: Additional results of structure and appearance control.    We present additional Ctrl-X results of structure and appearance control.
Additional structure and appearance control results.    We present additional results of structure and appearance control in Figure 14.
Figure 15: Appearance-only control.    Ctrl-X can do appearance-only control by dropping the structure control branch. Compared to IP-Adapter [43], our method shows better appearance alignment for both subjects and backgrounds.
Appearance-only control.    Ctrl-X is a method which disentangles control from given structure and appearance images, balancing structure alignment and appearance transfer when the two tasks are inherently conflicting. However, Ctrl-X can also achieve appearance-only control by simply dropping the structure control branch (and thus not needing to generate a structure image), as shown in Figure 15. Our method displays better appearance alignment for both subjects and background compared to the training-based IP-Adapter [43].
Figure 16: Structure-only control.    We display the jointly generated appearance images for prompt-driven conditional generation. Ctrl-X appearance transfer preserves the image quality of the generated appearances, so structure-only retains the quality of the base model.
Structure-only control.    For prompt-driven conditional (structure-only) generation, Ctrl-X needs to jointly generate an appearance image, where the jointly generated image is equivalent to vanilla SDXL v1.0 generation. We display the outputs alongside these appearance images in Figure 16, where there is minimal quality difference between the generated appearance images and the appearance-transferred output images, indicating that the need for appearance transfer does not greatly impact image quality. Thus, Ctrl-X adheres well to the quality of its base models.
@@ Appendix D Additional results
Figure 17: Extension to text-to-video (T2V) models.    Ctrl-X can be directly applied to T2V models for controllable video structure and appearance control, with AnimateDiff [9] with Realistic Vision v5.1 [32] and LaVie [39] here as examples. A playable video version of the AnimateDiff results can be found on our project page: https://genforce.github.io/ctrl-x/.
@@
Extension to video diffusion models.    We also present results of our method directly applied to text-to-video (T2V) diffusion models in Figure 17, namely AnimateDiff [9] with base model Realistic Vision v5.1 [32] and LaVie [39]. A playable video version of the AnimateDiff T2V results can be found on our project page: https://genforce.github.io/ctrl-x/.
Appendix E Dataset details
We publicly release our dataset in our code release: https://github.com/genforce/ctrl-x. In the release, we list all images present in the paper and their associated sources and licenses in a pdf file. All academic datasets which we use are cited here [3, 10, 43, 24, 48, 22, 19, 36, 16].
@@ Appendix E Dataset details
Overview.    Our dataset consists of 177 $1024\times 1024$ images divided into 16 types and across 7 categories. We split the images into condition images (67 images: “canny edge map”, “metadrive”, “3d mesh”, “3d humanoid”, “depth map”, “human pose image”, “point cloud”, “sketch”, “line drawing”, “HED edge drawing”, “normal map”, and “segmentation mask”) and natural images (110 images: “photo”, “painting”, “cartoon” and “birds eye view”), with the the largest type being “photo” (83 images). The condition images are further divided into two groups in our paper: ControlNet-supported conditions (“canny edge map”, “depth map”, “human pose image”, “line drawing”, “HED edge drawing”, “normal map”, and “segmentation mask”) and in-the-wild conditions (“metadrive”, “3D mesh”, “3D humanoid”, “point cloud”, and “sketch”). All of our images fall into one of seven categories: “animals” (52 images), “buildings” (11 images), “humans” (28 images), “objects” (29 images), “rooms” (24 images), “scenes” (22 images) and “vehicles” (11 images). About two thirds of the images come from the Web, while the remaining third is generated using SDXL 1.0 [27] or converted from natural images using Controlnet Annotators packaged in controlnet-aux [44]. For each of these images, we hand annotate them with a text prompt and other metadata (e.g. type). Then, these images, promtps, and metadata are combined to form the structure and appearance control dataset and conditional generation dataset, detailed below.
@@
In practice, when running Ctrl-X, we set the appearance prompt to be the same as the output prompt instead of our hand-annotated appearance prompt. We found little differences between the two.